OpenStudy (anonymous):

Let C be the curve of intersection of the parabolic cylinder x^2 = 2y and the surface 3z = xy. Find the exact length of C from the origin to the point (6,18,36)

OpenStudy (anonymous):

1. find the equation of line of intersection.

OpenStudy (anonymous):

6z=x^3

OpenStudy (amistre64):

a line integral is similar to arclength, except for a small modification .... right?

OpenStudy (amistre64):

pfft, line integral is something else ... :)

OpenStudy (amistre64):

is it easier to parametrise this, or .... other 3d coord it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

parametrization will reduce the number of integrals involved

OpenStudy (anonymous):

@amistre64 what substitution do you think?

OpenStudy (amistre64):

and the surfaces can be thought of as: z = x^2-2y z = xy/3

OpenStudy (anonymous):

x seems the highest order.. x = t?

OpenStudy (amistre64):

or, are you thinking of a y = x^2/2 sub :)

OpenStudy (anonymous):

i did that the first thing.. that is your "C"

OpenStudy (anonymous):

I did parameterize, but after getting \[\int\limits_{0}^{6} \sqrt(4 + 4t^2 + t^4)\] I don't know where the 1/2 comes from.

OpenStudy (anonymous):

x(t) = t y(t) = t^2/2 z(t) = t^3/6

OpenStudy (amistre64):

marvelous parametering ;)

OpenStudy (anonymous):

all hail distance formula

OpenStudy (anonymous):

@amistre64 that was your idea though

OpenStudy (anonymous):

Ya, then I plugged into the formula. But it has a 1/2 i believe... the normal distance formula doesn't have a 1/2..right?

OpenStudy (anonymous):

\[d(t)=\sqrt{x'^2+y'^2+z'^2}\]

OpenStudy (amistre64):

x(t) = t = 6 y(t) = t^2/2 = 18 z(t) = t^3/6 = 36 so from t=0, to t=6

OpenStudy (anonymous):

Still don't see where the 1/2 comes from...

OpenStudy (anonymous):

z'(t) = t^2/2

OpenStudy (amistre64):

x'(t) = 1 y'(t) = t z'(t) = t^2/2 \[\int_{0}^{6}\sqrt{1+t^2+\frac14t^4}~dt\]

OpenStudy (anonymous):

then how do you take that complex integral??

OpenStudy (amistre64):

\[\int_{0}^{6}\sqrt{1+t^2+\frac14t^4}~dt\] \[\int_{0}^{6}\sqrt{\frac14(4+4t^2+t^4)}~dt\] \[\frac12\int_{0}^{6}\sqrt{4+4t^2+t^4}~dt\] if need be

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

(2+t)^2

OpenStudy (anonymous):

i mean (2+t^2)^2

OpenStudy (anonymous):

ohhh I see where the 1/2 comes from. and then you just factor.

OpenStudy (amistre64):

soo\[\frac12\int_{0}^{6}t^2+2~dt\]doesnt seem so complex to me

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

haha, ya, didn't think to factor. Thanks!